evaluate the following integral
∫ 1 + cot x / x + log sin x dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello student,
Please find the answer to your question below

I = \int \frac{1+cotx}{x+log(sinx)}dx
x+log(sinx) = t
(1+\frac{cosx}{sinx})dx = dt
(1+cotx)dx = dt
I = \int \frac{dt}{t}
I = log(t) + c
I = log(x+log(sinx)) + c